Mark is 23 years older than his sister Ann. In 9 years, he will be twice as old as Ann. How old is each of them now?
Answer by mananth(16949)

Present age
Ann x
Mark' x + 23

After 9 years

Ann x + 9
Mark' ( x + 23 + 9 )
( x + 32 )

( x + 32 ) = 2 ( x + 9 )
1 x + 32 = 2 x + 18
x -2 x = 18 + -32
-1 x = -14

x= 14
Ann age = 14 years
Mark' age = 37 years